OpenStudy (anonymous):

m=8/9; (8,-9) equation ?

jimthompson5910 (jim_thompson5910):

\[\Large y = mx+b\] \[\Large -9 = \left(\frac{8}{9}\right)\left(8\right)+b\] \[\Large -9 = \frac{64}{9}+b\] \[\Large -9-\frac{64}{9} = b\] \[\Large \frac{-145}{9} = b\] \[\Large b = \frac{-145}{9}\] So the equation of the line is \[\Large y = \frac{8}{9}x-\frac{145}{9}\]